### v2

@@ 2.4 Parameter Learning and Loss Functions
The weights (parameters) of deep learning models are optimized using an optimization algorithm called gradient descent. However, it has to be noted that gradient descent is a generic algorithm which can be used to solve a wide range of optimization problems. In general, gradient descent finds the optimal weights by iteratively updating the weights such that the weights will result in a minimum prediction error over all instances in the training set. The prediction error is quantified by a loss function. For classification problems, the commonly used loss function is the negative log-likelihood loss or cross entropy loss, while the square loss and absolute loss are used for regression problems Wang et al. (2022). The weight update is defined as
$w_{i,j}^{l}=w_{i,j}^{l}-\alpha\nabla_{w_{i,j}}$
@@
The gradient can be computed across all instances of the training set, an approach known as batch gradient descent. However, this approach may not guarantee a convergence to the optimal solution as the gradient is the same for every weight update. An alternative approach is to perform the weight update on the basis of a single instance, but the approach results in a noisy gradient and becomes computationally intensive due to the frequent weight update. A more commonly used approach is to perform the weight update over a set of training instances, known as mini-batch gradient descent. This approach strikes a balance, providing a less noisy gradient and a more stable training process.
Several efforts have been made to improve the efficiency of gradient descent. One of the earlier efforts is the inclusion of past rate of change in the weight update to speed up the training of deep learning models, the algorithm is called gradient descent with momentum Qian (1999). Another effort is to improve the training convergence by adapting the learning rate based on the occurrence of the features Duchi et al. (2011). A more recent work utilizes both adaptive learning rate and momentum to improve the training efficiency and convergence of deep learning models Kingma and Ba (2017).
@@ 2.5 Regularization Methods
Regularization methods are employed to prevent overfitting in deep learning models and improve their generalization performance. Early stopping is a method that can detect the onset of overfitting during training by continuously monitoring the validation error. The model is considered overfitting if the validation error starts to increase at some point of the training while the training error is decreasing. However, detecting the onset of overfitting during the training of deep learning models is challenging due to the inherent stochasticity and the presence of noisy data. Several stopping criteria can be considered such as using a threshold to check if the decrease of (average) validation error is significant and count the number of successive increases of validation error Prechelt (2012).
Dropout is a regularization method that randomly switching off some neurons in the hidden layers during training with a predefined drop probability (dropout rate) Srivastava et al. (2014). Dropout has the effect of training and evaluating exponentially many different deep learning models. The dropout rate is a hyperparameter that needs to be carefully tuned to balance regularization and model capacity. Different ranges of dropout rate have been suggested. The original author suggested a dropout rate between 0.5 and 0.8 Srivastava et al. (2014) while others recommended a lower dropout rate between 0.1 and 0.3 Park and Kwak (2017). Also, it has been suggested a low dropout rate due to the exponential increase in the volume of training data Liu et al. (2023b).
Parameter norm penalty is a regularization method that adds a penalty term consisting of the network’s weights to the loss function. During the training, the penalty term discourages large weight values and hence, constraining the model’s capacity and reducing the chance of overfitting. The common penalty terms are $L^{1}$ norm penalty Tibshirani (1996), $L^{2}$ norm penalty, also known as weight decay and a combination of $L^{1}$ and $L^{2}$ Zou and Hastie (2005). An adaptive weight decay is proposed allowing the regularization strength for each weight to be dynamically adjusted Nakamura and Hong (2019).
Despite the advantages of the mini-batch gradient descent, each mini-batch may comprise data from different distributions. Furthermore, the data distribution may change after each weight update, which could slow down the training process. Batch normalization overcomes this issue by normalizing the summed input to a neuron over a mini batch of training instances Ioffe and Szegedy (2015). An alternative method is to perform normalization across the neurons instead of the mini batch, a method known as layer normalization Ba et al. (2016). Layer normalization is applicable in recurrent neural network and overcomes the dependencies on the mini batch size.
@@ 3.1 Deep Supervised Learning
Multilayer perceptron is a neural network model with one or more hidden fully-connected layers stacked between the input and output layers as shown in Fig. 6. The width (number of neurons) of the hidden layers and the depth (number of layers) of the network influence the model’s ability to learn patterns in the data. Specifically, the width determines the model’s ability to learn complex features while the depth allows the model to learn hierarchical representations of the data. Nevertheless, studies showed that a multilayer perceptron with a single hidden layer can approximate any continuous function Cybenko (1989), Hornik et al. (1989). Multilayer perceptron is effective in various industries and applications from healthcare to finance Widrow et al. (1994). However, multilayer perceptron requires the input data to be structured in a one-dimensional format e.g. tabular data, making it less suitable for unstructured data such as image, text and speech. To leverage multilayer perceptron for unstructured data, a feature extraction or transformation into structured data is necessary.
@@
Recurrent Neural Network is a neural network model that leverages the sequential information and memory through the use of recurrent connections, allowing it to effectively process data such as time series, text, speech and other sequential patterns. As shown in Fig. 7, a recurrent neural network is characterized by the recurrent connection which enables the network to loop back and use internal state from previous time step to the next time step. The internal state is parameterized by a set of weights which is shared across the sequence of the data. The training of recurrent neural networks suffers from the issue of vanishing gradient due to the challenges of propagation of gradients over a long sequence of data. Variants of recurrent neural networks are introduced to overcome the problem of vanishing gradient such as long short-term memory Hochreiter and Schmidhuber (1997) and gated recurrent memory Cho et al. (2014). The improved recurrent neural networks introduce memory cell and gating mechanisms to retain and discard information in every time step, allowing for more effective learning dependencies in long sequence. The network architecture can be built using fully-connected and convolutional layers Shi et al. (2015).
@@
AlexNet is among the first CNN models that gained widespread recognition and success, marking a significant achievement in the field of deep learning for computer vision tasks Krizhevsky et al. (2012). The model consists of five convolutional layers with max-pooling operation performed after the first and second convolutional layers, followed by three fully-connected layers. The first and second convolutional layers utilize a filter size of 11×11 and 5×5 respectively, and $3\times 3$ filter size is used for the remaining convolutional layers. ReLU activation function is used to mitigate the vanishing gradient.
@@
VGG-16 attempts to improve the CNN architecture by adding more convolutional layers, specifically up to 19 layers to capture more intricate feature representation from input data, followed by three fully-connected layers Simonyan and Zisserman (2015). ReLU activation function is used to reduce vanishing gradient. Unlike AlexNet, all convolutional layers utilize a small fix filter size of $3\times 3$ and max-pooling layer is added after a stack of two or three convolutional layers. This configuration allows the model to extract more discriminative features and decreases the number of parameters.
@@
ZFNet is a classic CNN model which has a similar architectural principle as AlexNet, featuring five convolutional layers with max-pooling layers after the first and second convolution, followed by three fully-connected layers Zeiler and Fergus (2013). The significant differences are the use of smaller filter size and stride in the convolutional layers and contrast normalization of the feature maps which allows the model to capture better features and improving the overall performance.
@@
Network-in-network introduces two innovative concepts to enhance the performance of the model Lin et al. (2014). The first was introducing a block of convolutional layers consisting of k×k convolution followed by two $1\times 1$ convolution operations. The pointwise convolutions are similar to applying multilayer perceptron on the feature maps, allowing the model to approximate more abstract feature representation. In the preceding models, the final feature maps are vectorized by flattening operation for classification by the fully-connected layers. Instead of flattening, network-in-network model calculates the spatial average of each feature map, and the resulting vector is fed to softmax function for classification. This approach is parameter-less, significantly reducing the number of parameters.
@@
GoogleNet leverages the fact that visual data can be represented at different scales by incorporating a module which consists of multiple convolutional pipelines with different filter sizes Szegedy et al. (2014). The module known as inception utilizes three kernel sizes ($5\times 5$, $3\times 3$, $1\times 1$) to capture spatial and channel information at different scales of resolution as shown in Fig. 9. This configuration enables a more effective feature extraction at both fine-grained and coarse-grained information from input data. The model architecture utilizes the inception module at the higher layers while the traditional convolution and max-pooling block is used to extract primitive and basic features. The inception modules are stacked upon each other with maximum pooling operation is performed occasionally to reduce the spatial resolution of the feature maps. GoogleNet utilizes global average pooling to vectorize the final feature maps before passing it to a fully-connected layer for classification.
@@
Increasing the number of layers enhances the model performance, mainly for solving complex tasks. However, training a very deep neural network is challenging due to the vanishing gradient problem, where the gradients that are used to update the network become insignificant or extremely small as they are backpropagated from the output layer to the earlier layers. A model called Highway Network overcomes this issue by introducing a gating mechanism that regulates the information flow of the layers, enabling the flow of information from the earlier layers to the later layers Srivastava et al. (2015). Consequently, this not only mitigates the vanishing gradient problem, but also renders the gradient-based training more tractable, enabling the training of very deep neural networks consisting as many as 100 layers.
The gating mechanism of Highway Network increases the number of parameters for regulating the information flow. ResNet is a CNN architecture that incorporates residual (skip) connection that allows information to bypass certain layers, mitigating the vanishing gradient problem He et al. (2015). ResNet architecture stacks residual blocks, which consists of two or three of convolutional layers with batch normalization and ReLU, and a skip connection which adds the input to the output of the final convolutional layer as shown in Fig. 10. If the input dimension does not match with the residual output dimension, a linear projection is performed by the residual connection to match the dimensions. In comparison to the gating mechanism of Highway Network, the residual connection is parameter-free, and thus does not incur additional computational costs. Furthermore, the connections are never closed whereby all information is always passed through the layers. This innovative concept enables the training of very deep neural networks boasting as many as 152 layers.
@@
DenseNet is another CNN architecture that overcomes the vanishing gradient problem. DenseNet follows the same approach as ResNet and Highway Network, utilizing skip connection to allow information flow from the earlier layers to later layers. However, DenseNet takes this concept one step further, by introducing a dense block consisting of multiple convolution functions (layers) with each convolution function performs batch normalization followed by ReLU and $3\times 3$ convolution. Each convolutional layer in the dense block receives feature maps from all its preceding layers, hence the connection is referred to as dense connection Huang et al. (2017). This configuration as shown in Fig. 11 maximizes information flow and preserves the feed-forward nature of the network. Dense blocks can become computationally expensive due to the increasing number of feature maps. To reduce the computational costs, a block of $1\times 1$ convolutional with batch normalization and max-pooling layers known as transition block is used to reduce the spatial dimension of the feature maps. The model architecture integrates these dense and transition blocks, stacking them alternately. The network depth can reach up to 264 layers.
@@
Although skip connections in ResNet effectively mitigate the vanishing gradient problem, a new challenge arises in the form of diminishing feature reuse as the network becomes deeper. Diminishing feature reuse refers to the diminishing effectiveness of the previously learned feature maps in subsequent layers, impacting the final prediction. WideResNet is a CNN architecture that is based on ResNet with the aim to mitigate diminishing feature reuse problem. Instead of making the network deeper, WideResNet makes the network wider by increasing the number of channels by k factor Zagoruyko and Komodakis (2017). The increased width allows the model to capture a more diverse features, enhancing its ability to learn complex relationships in the input data.
ResNext addresses the diminishing feature reuse by capturing more efficient and diverse features of the input data. ResNext introduces a concept of cardinality which is loosely based on the inception module as shown in Fig. 12. The cardinality refers to the number of independent and identical paths, where each path performs transformation of the input data, divided along the channel dimension Xie et al. (2017). In other words, instead of solely relying on increasing the depth of the model, ResNext enhances the feature learning by parallelizing the feature extraction through this cardinal path. In the proposed architecture, each path configuration is similar to the residual block of ResNet. The output from each path is then aggregated to form a comprehensive and diverse representation of the input data. The skip connection is used to mitigate the vanishing gradient problem. WideResNet is a CNN architecture that is based on ResNet with the aim to mitigate diminishing feature reuse problem. Instead of making the network deeper, WideResNet makes the network wider by increasing the number of channels by $k$ factor Zagoruyko and Komodakis (2017). The increased width allows the model to capture a more diverse features, enhancing its ability to learn complex relationships in the input data.
@@ 3.2 Deep Unsupervised Learning
Deep unsupervised models are trained with an unlabelled dataset. The learning process of these models involve calculating the prediction error through a loss function and utilizing the error to adjust the weights iteratively until the prediction error is minimized. Among the deep supervised models, three important models are identified namely multilayer perceptron, convolutional neural network and recurrent neural network.
Restricted Boltzmann Machine is a generative neural network model that learns a probability distribution based on a set of inputs. The model consists of a visible (input) layer and a hidden layer with symmetrically weighted connections as shown in Fig. 13. The input layer represents the input data with each node corresponding to a feature or variable while the hidden layer learns the abstract representation of the input data. Restricted Boltzmann machine model is trained using contrastive divergence, an algorithm that is based on a modified form of gradient descent, utilizing a sampling-based approach to estimate the gradient Hinton (2012). It has found success in solving combinative problems such as dimensionality reduction, collaborative filtering and topic modelling.
@@
Deep Belief Network can be viewed as a stack of restricted Boltzmann machines, comprising a visible layer and multiple hidden layers Hinton et al. (2006) as shown in Fig. 14. Deep belief network has two training phases. The initial phase is known as pretraining in which the network is trained layer by layer, with each layer serves as a pretraining layer of the subsequent layers. This sequential learning allows the hidden layers learn complex hierarchical feature representation of the data. The second phase is called fine-tuning whereby the deep belief network model can be further trained with supervision to perform tasks such as classification and regression Hinton (2009).
@@
Autoencoder is a generative neural network model that learns to encode the input data into a compressed representation and then reconstructs the original data from this representation. The layers that encode the input data is known as encoder while the layers that responsible for the reconstruction is referred to as the decoder as shown in Fig. 15. The encoded data (hidden layer) represents the abstract features of the input data also known as latent space or encoding. The decoder can be removed from the autoencoder, creating a standalone model that can be used for data compression and dimensionality reduction Romero et al. (2017), Li et al. (2020c). The decoder can also be replaced with predictive layers for classification task Mohd Noor (2021). The network architecture can be built using fully-connected and convolutional layers Li et al. (2023).
@@
Several autoencoder variants have been introduced to improve the autoencoder’s ability to capture better feature representation. Some introduced penalty terms to the loss function such as sparsity penalty (sparse autoencoder) Ng and others (2011) to encourage sparse representation and Jacobian Frobenius norm (contractive autoencoder) Rifai et al. (2011) to be less sensitive to small and insignificant variations in the input data while encoding the feature representation. Others trained the autoencoder to recover original data from corrupted data with noise Vincent et al. (2008). An improved denoising autoencoder knowns marginalized denoising autoencoder has been proposed which marginalizes the noise by adding a term that is linked to the encoding layer Chen et al. (2012). Variational autoencoder is a variant of autoencoder that has similar architecture i.e. encoder, latent space and decoder. Despite the similarity, instead of learning a fixed encoding, variational autoencoder learns the probability distribution of the input data in the latent space Kingma and Welling (2013). The model can be used to generate data by sampling from the learned probability distribution. The network architecture can be built by stacking more than one fully-connected layer and convolutional layer.
Generative Adversarial Network (GAN) is another generative neural network model that is designed for generating data that adheres closely to the distribution of the original training set. The model consists of two different neural networks namely generator and discriminator as shown in Fig. 16. The generator learns to imitate the distribution of the training set given a noise vector, effectively outsmarting the discriminator. Simultaneously during the training, the discriminator is trained to differentiate between the real data from the training set and synthetic data generated by the generator Goodfellow et al. (2014). This intricate dynamic between the networks drives an iterative learning process whereby the generator continually refines its ability to create synthetic data that closely resembles the real data while the discriminator enhances its ability to distinguish between authentic and fake data.
@@
The model can be extended by providing the labels to both generator and discriminator in which the model known as conditional GAN, capable of generating 1000 image classes Odena et al. (2017). Conditional GANs require a labelled dataset which might limit its application. InfoGAN is similar to conditional GAN, but the labels are substituted with latent codes, which allows the model to be trained in an unsupervised manner Chen et al. (2016). GANs often suffer from mode collapse whereby the model can only generate a single or small set of outputs. Wasserstein GAN improves the training by utilizing Wasserstein loss function which measures the difference between the real and synthesized data distribution Weng (2019). ProGAN tackles the training instability of GAN by progressively growing the generator and discriminator. The idea is that the model is scaled up gradually, starting with the simplest form of the problem and little by little the problem’s complexity is increased as the training progresses Karras et al. (2018). StyleGAN leverages the progressive GAN’s approach and neural style transfer to improve quality of the generated data Karras et al. (2019). The model is characterized by the independent manipulation of both style and content, allowing it to generate diverse styles and high quality data.

### v5

@@ 2.4 Parameter Learning and Loss Functions
The weights (parameters) of deep learning models are often optimized using an optimization algorithm called gradient descent, though other optimization algorithms may also be used. However, it has to be noted that gradient descent is a generic algorithm which can be used to solve a wide range of optimization problems. In general, gradient descent finds the optimal weights by iteratively updating the weights such that the weights will result in a minimum prediction error over all instances in the training set. The prediction error is quantified by a loss function. For classification problems, the commonly used loss function is the negative log-likelihood loss or cross entropy loss, while the square loss and absolute loss are used for regression problems [27]. The weights of layer $l$ is updated as
${w_{i,j}^{l}=w_{i,j}^{l}-\alpha\nabla_{w_{i,j}}}$ (8)
@@
The gradient can be computed across all training set instances, an approach known as batch gradient descent. However, this approach does not always guarantee convergence to the optimal solution, as it may get stuck in local minima or saddle points, and the same gradient is used for every weight update. An alternative approach is to perform the weight update on the basis of a single instance, but the approach results in a noisy gradient and becomes computationally intensive due to the frequent weight update. A more commonly used approach is to perform the weight update over a set of training instances, known as mini-batch gradient descent. This approach strikes a balance, providing a less noisy gradient and a more stable training process.
Several efforts have been made to improve the efficiency of gradient descent. One of the earlier efforts is the inclusion of past rate of change in the weight update to speed up the training of deep learning models, the algorithm is called gradient descent with momentum [28]. Another effort is to improve the training convergence by adapting the learning rate based on the occurrence of the features [29]. A more recent work utilizes both adaptive learning rate and momentum to improve the training efficiency and convergence of deep learning models [30].
@@ 2.5 Regularization Methods
Regularization methods are employed to prevent overfitting in deep learning models and improve their generalization performance. Early stopping is a method that can detect the onset of overfitting during training by continuously monitoring the validation error. The model is considered overfitting if the validation error starts to increase at some point of the training while the training error is decreasing. However, detecting the onset of overfitting during the training of deep learning models is challenging due to the inherent stochasticity and the presence of noisy data. Several stopping criteria can be considered such as using a threshold to check if the decrease of (average) validation error is significant and count the number of successive increases of validation error [31].
Dropout is a regularization method that randomly switching off some neurons in the hidden layers during training with a predefined drop probability (dropout rate) [32]. Dropout has the effect of training and evaluating exponentially many different deep learning models. The dropout rate is a hyperparameter that needs to be carefully tuned to balance regularization and model capacity. Different ranges of dropout rate have been suggested. The original author suggested a dropout rate between 0.5 and 0.8 [32] while others recommended a lower dropout rate between 0.1 and 0.3 [33]. Also, it has been suggested a low dropout rate due to the exponential increase in the volume of training data [34].
Parameter norm penalty is a regularization method that adds a penalty term consisting of the network’s weights to the loss function. During the training, the penalty term discourages large weight values and hence, constraining the model’s capacity and reducing the chance of overfitting. The common penalty terms are $L^{1}$ norm penalty [35], $L^{2}$ norm penalty, also known as weight decay and a combination of $L^{1}$ and $L^{2}$ [36]. An adaptive weight decay is proposed allowing the regularization strength for each weight to be dynamically adjusted [37].
Despite the advantages of the mini-batch gradient descent, each mini-batch may comprise data from different distributions. Furthermore, the data distribution may change after each weight update, which could slow down the training process. Batch normalization overcomes this issue by normalizing the summed input to a neuron over a mini batch of training instances [38]. An alternative method is to perform normalization across the neurons instead of the mini batch, a method known as layer normalization [39]. Layer normalization is applicable in recurrent neural network and overcomes the dependencies on the mini batch size.
@@ 3.1 Deep Supervised Learning
Multilayer perceptron is a neural network model with one or more hidden fully-connected layers stacked between the input and output layers as shown in Fig. 7. The width (number of neurons) of the hidden layers and the depth (number of layers) of the network influence the model’s ability to learn patterns in the data. Specifically, the width affects the network’s ability to capture a broader range of features while the depth facilitates the learning of hierarchical representations. Nevertheless, studies showed that a multilayer perceptron with a single hidden layer can approximate any continuous function [40], [41]. Multilayer perceptron is effective in various industries and applications from healthcare to finance [42]. However, multilayer perceptron requires the input data to be structured in a one-dimensional format e.g. tabular data, making it less suitable for unstructured data such as image, text and speech. To leverage multilayer perceptron for unstructured data, a feature extraction or transformation into structured data is necessary.
@@
Recurrent Neural Network (RNN) is a neural network model that leverages the sequential information and memory through the use of recurrent connections, allowing it to effectively process data such as time series, text, speech and other sequential patterns. As shown in Fig. 8, a recurrent neural network is characterized by the recurrent connection which enables the network to loop back and use internal state from the previous time step to the next time step. The internal state is parameterized by a set of weights shared across the sequence of data. The training of recurrent neural networks suffers from the issue of vanishing gradient due to the challenges of propagation of gradients over a long sequence of data. Variants of recurrent neural networks are introduced to overcome the problem of vanishing gradient such as long short-term memory (LSTM) [43] and gated recurrent memory (GRU) [44]. The improved recurrent neural networks introduce memory cell and gating mechanisms to retain and discard information in every time step, allowing for more effective learning dependencies in long sequence. The network architecture can be built using fully-connected and convolutional layers [45].
@@
AlexNet is among the first CNN models that gained widespread recognition and success, marking a significant achievement in the field of deep learning for computer vision tasks [46]. The model consists of five convolutional layers with max-pooling operation performed after the first and second convolutional layers, followed by three fully-connected layers. The first and second convolutional layers utilize a filter size of 11×11 and 5×5 respectively, and $3\times 3$ filter size is used for the remaining convolutional layers. ReLU activation function is used to mitigate the vanishing gradient.
@@
VGG-16 attempts to improve the CNN architecture by adding more convolutional layers, specifically up to 19 layers to capture more intricate feature representation from input data, followed by three fully-connected layers [47]. ReLU activation function is used to reduce vanishing gradient. Unlike AlexNet, all convolutional layers utilize a small fix filter size of $3\times 3$ and max-pooling layer is added after a stack of two or three convolutional layers. This configuration allows the model to extract more discriminative features and decreases the number of parameters.
@@
ZFNet is a classic CNN model which has a similar architectural principle as AlexNet, featuring five convolutional layers with max-pooling layers after the first and second convolution, followed by three fully-connected layers [48]. The significant differences are the use of smaller filter size and stride in the convolutional layers and contrast normalization of the feature maps which allows the model to capture better features and improving the overall performance.
@@
Network-in-network introduces two innovative concepts to enhance the performance of the model [49]. The first was introducing a block of convolutional layers consisting of k×k convolution followed by two $1\times 1$ convolution operations. The pointwise convolutions are similar to applying multilayer perceptron on the feature maps, allowing the model to approximate more abstract feature representation. In the preceding models, the final feature maps are vectorized by flattening operation for classification by the fully-connected layers. Instead of flattening, network-in-network model calculates the spatial average of each feature map, and the resulting vector is fed to softmax function for classification. This approach is parameter-less, significantly reducing the number of parameters.
@@
GoogleNet leverages the fact that visual data can be represented at different scales by incorporating a module which consists of multiple convolutional pipelines with different filter sizes [50]. The module known as inception utilizes three kernel sizes ($5\times 5$, $3\times 3$, $1\times 1$) to capture spatial and channel information at different scales of resolution as shown in Fig. 10. This configuration enables a more effective feature extraction at both fine-grained and coarse-grained information from input data. The model architecture utilizes the inception module at the higher layers while the traditional convolution and max-pooling block is used to extract primitive and basic features. The inception modules are stacked upon each other with maximum pooling operation is performed occasionally to reduce the spatial resolution of the feature maps. GoogleNet utilizes global average pooling to vectorize the final feature maps before passing it to a fully-connected layer for classification.
@@
Increasing the number of layers enhances the model performance, mainly for solving complex tasks. However, training a very deep neural network is challenging due to the vanishing gradient problem, where the gradients that are used to update the network become insignificant or extremely small as they are backpropagated from the output layer to the earlier layers. A model called Highway Network overcomes this issue by introducing a gating mechanism that regulates the information flow of the layers, enabling the flow of information from the earlier layers to the later layers [51]. Consequently, this not only mitigates the vanishing gradient problem, but also renders the gradient-based training more tractable, enabling the training of very deep neural networks consisting as many as 100 layers.
The gating mechanism of Highway Network increases the number of parameters for regulating the information flow. ResNet is a CNN architecture that incorporates residual (skip) connection that allows information to bypass certain layers, mitigating the vanishing gradient problem [52]. ResNet architecture stacks residual blocks, which consists of two or three of convolutional layers with batch normalization and ReLU, and a skip connection which adds the input to the output of the final convolutional layer as shown in Fig. 11. If the input dimension does not match with the residual output dimension, a linear projection is performed by the residual connection to match the dimensions. In comparison to the gating mechanism of Highway Network, the residual connection is parameter-free, and thus does not incur additional computational costs. Furthermore, the connections are never closed whereby all information is always passed through the layers. This innovative concept enables the training of very deep neural networks boasting as many as 152 layers.
@@
DenseNet is another CNN architecture that overcomes the vanishing gradient problem. DenseNet follows the same approach as ResNet and Highway Network, utilizing skip connection to allow information flow from the earlier layers to later layers. However, DenseNet takes this concept one step further, by introducing a dense block consisting of multiple convolution functions (layers) with each convolution function performs batch normalization followed by ReLU and $3\times 3$ convolution. Each convolutional layer in the dense block receives feature maps from all its preceding layers, hence the connection is referred to as dense connection [53]. This configuration as shown in Fig. 12 maximizes information flow and preserves the feed-forward nature of the network. Dense blocks can become computationally expensive due to the increasing number of feature maps. To reduce the computational costs, a block of $1\times 1$ convolutional with batch normalization and max-pooling layers known as transition block is used to reduce the spatial dimension of the feature maps. The model architecture integrates these dense and transition blocks, stacking them alternately. The network depth can reach up to 264 layers.
@@
Although skip connections in ResNet effectively mitigate the vanishing gradient problem, a new challenge arises in the form of diminishing feature reuse as the network becomes deeper. Diminishing feature reuse refers to the diminishing effectiveness of the previously learned feature maps in subsequent layers, impacting the final prediction. WideResNet is a CNN architecture that is based on ResNet with the aim to mitigate diminishing feature reuse problem. Instead of making the network deeper, WideResNet makes the network wider by increasing the number of channels by k factor [54]. The increased width allows the model to capture a more diverse features, enhancing its ability to learn complex relationships in the input data.
ResNext addresses the diminishing feature reuse by capturing more efficient and diverse features of the input data. ResNext introduces a concept of cardinality which is loosely based on the inception module as shown in Fig. 13. The cardinality refers to the number of independent and identical paths, where each path performs transformation of the input data, divided along the channel dimension [55]. In other words, instead of solely relying on increasing the depth of the model, ResNext enhances the feature learning by parallelizing the feature extraction through this cardinal path. In the proposed architecture, each path configuration is similar to the residual block of ResNet. The output from each path is then aggregated to form a comprehensive and diverse representation of the input data. The skip connection is used to mitigate the vanishing gradient problem. WideResNet is a CNN architecture that is based on ResNet with the aim to mitigate diminishing feature reuse problem. Instead of making the network deeper, WideResNet makes the network wider by increasing the number of channels by $k$ factor [54]. The increased width allows the model to capture a more diverse features, enhancing its ability to learn complex relationships in the input data.
@@ 3.2 Deep Unsupervised Learning
Deep unsupervised models are trained with an unlabelled dataset. The learning process of these models discovering patterns, structures and representations within the data without relying on explicit labels or supervision. Instead, these models often learn by optimizing objective functions that capture the underlying data characteristics such as clustering, learning useful feature embeddings and reconstructing input data from compressed representations. Examples of deep unsupervised models are autoencoders, generative adversarial networks and restricted Boltzmann machines.
Restricted Boltzmann Machine is a generative neural network model that learns a probability distribution based on a set of inputs. The model consists of a visible (input) layer and a hidden layer with symmetrically weighted connections as shown in Fig. 14. The input layer represents the input data with each node corresponding to a feature or variable while the hidden layer learns the abstract representation of the input data. Restricted Boltzmann machine model is trained using contrastive divergence, an algorithm that is based on a modified form of gradient descent, utilizing a sampling-based approach to estimate the gradient [56]. It has found success in solving combinative problems such as dimensionality reduction, collaborative filtering and topic modelling.
@@
Deep Belief Network can be viewed as a stack of restricted Boltzmann machines, comprising a visible layer and multiple hidden layers [57] as shown in Fig. 15. Deep belief network has two training phases. The initial phase is known as pretraining in which the network is trained layer by layer, with each layer serves as a pretraining layer of the subsequent layers. This sequential learning allows the hidden layers learn complex hierarchical feature representation of the data. The second phase is called fine-tuning whereby the deep belief network model can be further trained with supervision to perform tasks such as classification and regression [58].
@@
Autoencoder is a generative neural network model that learns to encode the input data into a compressed representation and then reconstructs the original data from this representation. The layers that encode the input data is known as encoder while the layers that responsible for the reconstruction is referred to as the decoder as shown in Fig. 16. The encoded data (hidden layer) represents the abstract features of the input data also known as latent space or encoding. The decoder can be removed from the autoencoder, creating a standalone model that can be used for data compression and dimensionality reduction [59], [60]. The decoder can also be replaced with predictive layers for classification task [61]. The network architecture can be built using fully-connected and convolutional layers [62].
@@
Several autoencoder variants have been introduced to improve the autoencoder’s ability to capture better feature representation. Some introduced penalty terms to the loss function such as sparsity penalty (sparse autoencoder) [63] to encourage sparse representation and Jacobian Frobenius norm (contractive autoencoder) [64] to be less sensitive to small and insignificant variations in the input data while encoding the feature representation. Others trained the autoencoder to recover original data from corrupted data with noise [65]. An improved denoising autoencoder knowns marginalized denoising autoencoder has been proposed which marginalizes the noise by adding a term that is linked to the encoding layer [66]. Variational autoencoder is a variant of autoencoder that has similar architecture i.e. encoder, latent space and decoder. Despite the similarity, instead of learning a fixed encoding, variational autoencoder learns the probability distribution of the input data in the latent space [67]. The model can be used to generate data by sampling from the learned probability distribution. The network architecture can be built by stacking more than one fully-connected layer and convolutional layer.
Generative Adversarial Network (GAN) is another generative neural network model that is designed for generating data that adheres closely to the distribution of the original training set. The model consists of two different neural networks namely generator and discriminator as shown in Fig. 17. The generator learns to imitate the distribution of the training set given a noise vector, effectively outsmarting the discriminator. Simultaneously during the training, the discriminator is trained to differentiate between the real data from the training set and synthetic data generated by the generator [68]. This intricate dynamic between the networks drives an iterative learning process whereby the generator continually refines its ability to create synthetic data that closely resembles the real data while the discriminator enhances its ability to distinguish between authentic and fake data.
@@
The model can be extended by providing the labels to both generator and discriminator in which the model known as conditional GAN, capable of generating 1000 image classes [69]. Conditional GANs require a labelled dataset which might limit its application. InfoGAN is similar to conditional GAN, but the labels are substituted with latent codes, which allows the model to be trained in an unsupervised manner [70]. GANs often suffer from mode collapse whereby the model can only generate a single or small set of outputs. Wasserstein GAN improves the training by utilizing Wasserstein loss function which measures the difference between the real and synthesized data distribution [71]. ProGAN tackles the training instability of GAN by progressively growing the generator and discriminator. The idea is that the model is scaled up gradually, starting with the simplest form of the problem and little by little the problem’s complexity is increased as the training progresses [72]. StyleGAN leverages the progressive GAN’s approach and neural style transfer to improve quality of the generated data [73]. The model is characterized by the independent manipulation of both style and content, allowing it to generate diverse styles and high quality data.